# Quantum mixing on large Schreier graphs

### v1

@@ Abstract
Quantum ergodicity describes the delocalization of most eigenfunctions of Laplace-type operators on graphs or manifolds exhibiting chaotic classical dynamics. Quantum mixing is a stronger notion, additionally controlling correlations between eigenfunctions at different energy levels. In this work, we study families of finite Schreier graphs that converge to an infinite Cayley graph and establish quantum mixing under the assumption that the limiting Cayley graph has absolutely continuous spectrum. The convergence of Schreier graphs is understood in the Benjamini–Schramm sense or in the sense of strong convergence in distribution.
Our proofs rely on a new approach to quantum ergodicity, based on trace computations, resolvent approximations and representation theory. We illustrate our assumptions on several examples and provide applications to Schreier graphs associated with free products of groups and right-angled Coxeter groups.
@@ 1 Introduction and main results
Eigenvalue statistics of quantum systems are expected to reflect the dynamical properties of their classical counterparts. Berry and Tabor conjectured that for integrable systems, eigenvalues tend to behave “universally” like independent random variables (Poisson statistics). In contrast, for chaotic systems, Bohigas, Giannoni and Schmit conjectured that eigenvalues tend to follow random matrix theory statistics (GOE, GUE, etc.).
@@
Beyond regular graphs, the paper [6] proved that quantum ergodicity holds along sequences of not necessarily regular graphs. The result is established under some assumptions on the spectrum of a limiting object. This limiting object is the Benjamini-Schramm limit of the sequence of graphs. One of the assumptions is a bound on the resolvent of its underlying Schrödinger operator which implies absolute continuity of the spectrum, and it has been verified in several concrete examples in [5, 7]. In a nutshell, [6] establishes in a rather general setting that “spectral delocalization implies spatial delocalization of eigenvectors”. As in the original works on quantum ergodicity on manifolds [67, 81, 31, 68], the proofs are based on propagation arguments using a classical flow, here the non-backtracking flow on the graph. The results of [6, 5, 7] are however restricted to large discrete graphs which look like trees near most points (i.e., the Benjamini-Schramm limit is supported on infinite trees). Beyond this tree-like setting, quantum ergodicity was proved in [62] for the family of $\mathbb{Z}^{d}$-periodic graphs, away from flat bands (point spectrum). This relied in an essential way on Bloch-Floquet theory.
@@ 1.1 Setup
It is equipped with the involution $p^{*}=\sum_{g}\bar{p}_{g}g^{-1}$. If $p\in\mathbb{C}[\Gamma]$ and $\rho$ is a unitary representation of $\Gamma$ on a Hilbert space $H$, we define the operator in $\mathcal{B}(H)$
@@
We first define small-scale quantum weak mixing and postpone to Section 1.7 a tentative definition of quantum weak mixing in this framework, as well as some discussions on related works in random matrix theory. We start with the simplest case $a_{N}\in\mathbb{C}^{N}$ with $\|a_{N}\|_{\infty}\leq 1$. We shall say that small-scale quantum weak mixing occurs if for all $E_{1},E_{2}$ in the interior of $I$,
@@
where $J_{E}^{\eta}=[E-\eta,E+\eta]$ (notice that for all $\beta\neq\alpha$, $\langle\varphi_{\beta},a_{N}\varphi_{\alpha}\rangle=\langle\varphi_{\beta},a_{N}\varphi_{\alpha}\rangle-\langle a_{N}\rangle\langle\varphi_{\beta},\varphi_{\alpha}\rangle$). Compared to (1.8), (1.10) also captures some information on the correlations between distinct eigenvectors. For simplicity, we will often speak of quantum mixing in place of small-scale quantum weak mixing from now on.
@@
As before, there is a generalization of quantum mixing to non-necessarily diagonal observables. We fix a finite subset $T\subset\Gamma$ and consider a sequence $K_{N}\in M_{N}(\mathbb{C})$ of $T$-local matrices as in (1.7) and such that $\|K_{N}\|_{1,\infty}\leq 1$ for any $N$. We say that small-scale quantum weak mixing (or simply quantum mixing) occurs if for any $E_{1},E_{2}$ in the interior of $I$,
@@ 1.2 Quantum mixing for asymptotically uncorrelated observables
Assume that $\rho_{N}$ converges in distribution toward $\lambda$. Let $p\in\mathbb{C}[\Gamma]$ such that $p=p^{*}$ and an interval $I_{0}$ such that (1.5) holds for $P=\lambda(p)$. Let $I$ be a closed interval in the interior of $I_{0}$ and $(\varphi_{\alpha})$ be an orthonormal eigenbasis $P_{N}=\rho_{N}(p)$. Then, for any finite $T\subset\Gamma$ and any asymptotically uncorrelated sequence of $T$-local matrices $(K_{N})$ in $M_{N}(\mathbb{C})$ such that $\|K_{N}\|_{1,\infty}\leq 1$, quantum ergodicity (1.8) and quantum mixing (1.11) hold in $I$.
@@ 1.7 Further discussion and open questions
Quantum weak mixing. We have defined small-scale quantum mixing in (1.10)-(1.11). Such small-scale estimates are very natural from the perspective of local laws established in random matrix theory. In analogy with [83, 84], one may also be interested in quantum weak mixing in a macroscopic interval $I$. A reasonable definition for this would be that, for the sequence of $T$-local observables $(K_{N})$, and for any real $\tau$, we have
@@
On the lower bound on the resolvent in (1.5). The lower bound $\Im R^{z}(e,e)\geq C_{0}^{-1}$ in (1.5) is only needed in the proofs of small scale quantum mixing to guarantee that for some $C^{\prime}_{0}>0$, for any $\eta>0$, all $N$ large enough and all $E\in I$, we have $|\Lambda_{J^{\eta}_{E}}|\geq C^{\prime}_{0}N\eta$. If one is interested in quantum ergodicity (1.8) or quantum weak mixing (1.18) on a fixed interval $I\subset I_{0}$, then a lower bound of the form $|\Lambda_{I}|\geq C^{\prime}_{0}N$ is sufficient (where $C^{\prime}_{0}$ is allowed to depend on $I$). The latter condition is satisfied under the sole hypothesis that the spectrum of the operator $P=\lambda(p)$ intersects the interior of $I$. See Remark 3.3 for more details.
@@ 2 Quantum mixing and resolvent estimates
2.1 Small scale quantum mixing implies quantum ergodicity and quantum weak mixing
@@
with the convention that $L_{I,J}=L_{I}^{\tau,\eta}=0$ if $\Lambda_{I}$ is empty. The notation is slightly abusive since $L^{\tau,\eta}_{I}(A,K)$ actually depends on the choice of the orthonormal basis. The quantum moment $L_{IJ}(A,K)$ is however independent of the choice of the basis as it can be written in terms of spectral projections (see the proof of Lemma 2.2 below). Note that, for centered observables $K$, the term $L^{0,0}_{I}$ appears in the definition of quantum ergodicity (1.8), the term $L^{\tau,\eta}_{I}$ in the definition of quantum weak mixing (1.18), while $L_{IJ}$ appears in the definition of small-scale quantum weak mixing (1.11). We will be interested in situations where these moments are small. Since $L_{IJ}(A,K)$ is unitarily invariant, it will be easier to upper bound. However for $|\Lambda_{J}|$ large, $L_{IJ}(A,K)$ cannot be expected to be small in general.
Our strategy to upper bound $L_{I}^{\tau,\eta}$ will be to decompose $I$ into small sets counting few eigenvalues. The following lemma shows that small scale quantum mixing implies quantum ergodicity and quantum weak mixing (using the second part of the statement with $K_{N}-\langle K_{N}\rangle$ instead of $K_{N}$).
Lemma 2.1 (Small scale quantum mixing implies quantum ergodicity and mixing).
@@ 2.2 A general bound on quantum variance
The quantity $L_{IJ}(A,K)$ introduced in (2.1) is the central object that must be bounded from above in order to prove small-scale quantum mixing. One key tool to control this quantity is to observe that it can be expressed as a trace involving $K$ and two functions of the operator $A$.
@@ 2.3 Proof ideas
$\left|\{g\in\Gamma\mid gx=y\text{ and }f(p)_{g}\neq 0\}\right|>1.$
@@
$|f(P_{N})(x,y)|^{2}=\Bigl|\sum_{g\in\Gamma,\ g\cdot x=y}f(p)_{g}\Bigr|^{2}=\sum_{g\in\Gamma,\ g\cdot x=y}|f(p)_{g}|^{2}=\rho_{N}(q)(x,y)$ (2.6)
@@ 2.5 Resolvent identities
We explained in the previous section that the imaginary part of the resolvent appears in the trace computations needed to prove our main results. We gather here some elementary facts regarding resolvents. Let $X$ be a countable set and $P$ be a bounded self-adjoint operator on $\ell^{2}(X)$. In our case, we are interested in $X=\Gamma$, $P=\lambda(p)$ with $p=p^{*}\in\mathbb{C}[\Gamma]$. For $z\in\mathbb{C}^{+}=\{z\in\mathbb{C}:\Im(z)>0\}$, we define the resolvent of $P$ as $R^{z}=(P-z\mathrm{Id})^{-1}$. We will make use of the following resolvent identity, known as Ward’s identity: with $\eta=\Im(z)$,
@@
where $x\in X$. It is a consequence of the fact that $\sum_{y\in X}|R^{z}(x,y)|^{2}=(R^{z}(R^{z})^{*})(x,x)$ together with the spectral theorem which implies that $\eta R^{z}(R^{z})^{*}=\Im R$ since $\Im(1/z)=-\Im(z)/|z|^{2}$ and $\Im(P-z\mathrm{Id})=-\eta\mathrm{Id}$, $|P-z\mathrm{Id}|^{-2}=R^{z}(R^{z})^{*}$. In the same vein, writing $z=E+i\eta$ and denoting by $\mu_{P}^{\delta_{x}}$ the spectral measure at vector $\delta_{x}$ (see (3.6)), we have
@@ 3.2 The number of eigenvalues in small intervals
We next take as usual $\rho_{N}$ converging in distribution toward $\lambda$ as $N\to\infty$. For an interval $J\subset I$, our goal in this subsection is to compare the number of eigenvalues of $\rho_{N}(p)$ in $J$ and $N\mu_{p}(J)$. For this, we prove the weak convergence to $\mu_{p}$ of the empirical distribution of eigenvalues of $\rho_{N}(p)$. More precisely, we establish the following quantitative lemma.
@@ 3.3 Trace estimate for local operators
$A(x,y)=\sum_{g\in T}a_{g}(x){\mathbf{1}}(x=g.y).$ (3.11)
Such a decomposition always exists: if for fixed $x,y$ there is a unique $g$ such that $x=g.y$, then $a_{g}(x)$ is uniquely defined by $a_{g}(x):=A(x,g^{-1}.x)$. This will be the typical situation considered below. If not, the decomposition is not unique. One could for example pick arbitrarily one $g$ such that $x=g.y$, put $a_{g}(x)=A(x,g^{-1}.x)$ and $a_{t}(x)=0$ for $t\neq g$ such that $t^{-1}.x=g^{-1}.x=y$. In the sequel, we make the more canonical choice $a_{g}(x)=A(x,g^{-1}.x)/m_{g}(x)$ where $m_{g}(x):=\sum_{t\in T}\mathbf{1}(g^{-1}.x=t^{-1}.x)\geq 1$.
@@
The goal of this section is to prove Corollary 3.5, which will be used in conjunction with Lemma 2.2. Since we are willing to bound traces of the form ${\rm Tr}(K_{1}P_{1}K_{2}^{*}P_{2})$, we express them in a simpler form where $P_{1}$ and $P_{2}$ are merged into a single local operator $Q$, up to a remainder term controlled thanks to the assumption of convergence in distribution. In the case of diagonal observables $K_{1},K_{2}$ (corresponding to $T_{1}=T_{2}=\{e\}$ below), the statement simplifies dramatically and boils down to (2.5)-(2.6)-(2.7).
@@ 4.1 Quantum mixing estimate for Schreier graphs
We start by choosing $f_{j}$ and give some of its properties. We set
$\epsilon=\frac{1}{\ln\left(\max\left(\frac{1}{\eta},4\right)\right)}$ (4.3)
@@
$\leq 4\sqrt{\int\frac{\eta^{2}}{(\lambda-E_{j})^{2}+\eta^{2}}d\mu_{p}(\lambda)}+\min(4\eta,1)$
@@
We may now evaluate (4.2) for our choice of $f_{j}$. We denote by $S$ a finite symmetric generating set of $\Gamma$ containing the support of $p$ and $t$. For integer $r\geq 1$, $B_{S}(r)$ denotes the ball of radius $r$ in $\mathrm{Cay}(\Gamma,S)$ and ${\rm Bad}_{S}(r)=\mathrm{Fix}^{*}_{N}(B_{S}(2r))$. From Corollary 3.5, we have
@@
Since $f_{i}\geq 0$, we have $q=a^{*}a$ for some $a\in\mathbb{C}[\Gamma]$ as a consequence of Schur product theorem in $\mathbb{C}[\Gamma]$ [^1] See this post on mathoverflow. and thus $Q_{N}=\rho_{N}(q)=\rho_{N}(a)^{*}\rho_{N}(a)$ is non-negative. From (4.4)-(4.5) and $\|k_{N}\|_{\infty}\leq 1$, we deduce our main upper bound: for some new $C>0$,
@@
Depending on the various assumptions stated in Theorems 1.3, 1.5 and 1.7, we will prove upper bounds on
$\frac{\langle k_{N},Q_{N}k_{N}\rangle}{N\eta}$
which will imply (4.9).
@@
where we have used the bound $|q_{g}|=|f_{1}(p)|_{t^{-1}gt}|f_{2}(p)_{g}|\leq C_{1}^{2}\eta^{2}$ from (4.6). Note that this upper bound is uniform over all $E_{1},E_{2}\in I$. We deduce that
@@
Since $\delta$ can be arbitrary small, this proves (4.9) and concludes the proof of Theorem 1.3.∎

### v2

@@ Abstract
We prove quantum ergodicity and quantum mixing for sequences of finite Schreier graphs converging to an infinite Cayley graph whose adjacency operator has absolutely continuous spectrum. Under Benjamini–Schramm convergence (or strong convergence in distribution), we show that correlations between eigenvectors at distinct energies vanish asymptotically when tested against a broad class of local observables.
Our results apply to all orthonormal eigenbases and do not require tree-like structure or periodicity of the limiting graph, unlike previous approaches based on non-backtracking operators or Floquet theory. The proof introduces a new framework for quantum ergodicity, based on trace identities, resolvent approximations and representation-theoretic techniques and extends to certain families of non-regular graphs.
We illustrate the assumptions and consequences of our theorems on Schreier graphs arising from free products of groups, right-angled Coxeter groups and lifts of a fixed base graph.
@@ 1 Introduction and main results
Eigenvalue statistics of quantum systems are expected to reflect the dynamical properties of their classical counterparts. Berry and Tabor conjectured that for integrable systems, eigenvalues tend to behave universally like independent random variables (Poisson statistics). In contrast, for chaotic systems, Bohigas, Giannoni and Schmit conjectured that eigenvalues tend to follow random matrix theory statistics (GOE, GUE, etc.).
@@
Beyond regular graphs, the paper [6] proved that quantum ergodicity holds along sequences of not necessarily regular graphs. The result is established under some assumptions on the spectrum of a limiting object. This limiting object is the Benjamini-Schramm limit of the sequence of graphs. One of the assumptions is a bound on the resolvent of its underlying Schrödinger operator which implies absolute continuity of the spectrum, and it has been verified in several concrete examples in [5, 7]. In a nutshell, [6] establishes in a rather general setting that “spectral delocalization implies spatial delocalization of eigenvectors”. As in the original works on quantum ergodicity on manifolds [67, 81, 31, 68], the proofs are based on propagation arguments using a classical flow, here the non-backtracking flow on the graph. The results of [6, 5, 7], however, are restricted to large discrete graphs which look like trees near most points (i.e., the Benjamini-Schramm limit is supported on infinite trees). Beyond this tree-like setting, quantum ergodicity was proved in [62] for the family of $\mathbb{Z}^{d}$-periodic graphs, away from flat bands (point spectrum). This relied in an essential way on Bloch-Floquet theory.
@@ 1.1 Setup
It is equipped with the involution $p^{*}=\sum_{g}\bar{p}_{g}g^{-1}$ and the multiplication is $pq=\sum_{g,h}p_{g}q_{h}\,gh$. If $p\in\mathbb{C}[\Gamma]$ and $\rho$ is a unitary representation of $\Gamma$ on a Hilbert space $H$, we define the operator in $\mathcal{B}(H)$
@@
We first define small scale quantum weak mixing and postpone to Section 1.7 a tentative definition of quantum weak mixing in this framework, as well as some discussions on related works in random matrix theory. We start with the simplest case $a_{N}\in\mathbb{C}^{N}$ with $\|a_{N}\|_{\infty}\leq 1$. We shall say that small scale quantum weak mixing occurs if for all $E_{1},E_{2}$ in the interior of $I$,
@@
where $J_{E}^{\eta}=[E-\eta,E+\eta]$ (notice that for all $\beta\neq\alpha$, $\langle\varphi_{\beta},a_{N}\varphi_{\alpha}\rangle=\langle\varphi_{\beta},a_{N}\varphi_{\alpha}\rangle-\langle a_{N}\rangle\langle\varphi_{\beta},\varphi_{\alpha}\rangle$). Compared to (1.8), (1.10) also captures some information on the correlations between distinct eigenvectors. For simplicity, we will often speak of quantum mixing in place of small scale quantum weak mixing from now on.
@@
As before, there is a generalization of quantum mixing to non-necessarily diagonal observables. We fix a finite subset $T\subset\Gamma$ and consider a sequence $K_{N}\in M_{N}(\mathbb{C})$ of $T$-local matrices as in (1.7) and such that $\|K_{N}\|_{1,\infty}\leq 1$ for any $N$. We say that small scale quantum weak mixing (or simply quantum mixing) occurs if uniformly over $E_{1},E_{2}$ in the interior of $I$,
@@ 1.2 Quantum mixing for asymptotically uncorrelated observables
Assume that $\rho_{N}$ converges in distribution toward $\lambda$. Let $p\in\mathbb{C}[\Gamma]$ such that $p=p^{*}$ and an interval $I_{0}$ such that (1.5) holds for $P=\lambda(p)$. Let $I$ be a closed interval in the interior of $I_{0}$ and $(\varphi_{\alpha})$ be an orthonormal eigenbasis for $P_{N}=\rho_{N}(p)$. Then, for any finite $T\subset\Gamma$ and any asymptotically uncorrelated sequence of $T$-local matrices $(K_{N})$ in $M_{N}(\mathbb{C})$ such that $\|K_{N}\|_{1,\infty}\leq 1$, quantum ergodicity (1.8) and quantum mixing (1.11) hold in $I$.
@@ 1.7 Further discussion and open questions
Quantum weak mixing. We have defined small scale quantum weak mixing in (1.10)-(1.11). Such small scale estimates are very natural from the perspective of local laws established in random matrix theory. In analogy with [83, 84], one may also be interested in quantum weak mixing in a macroscopic interval $I$. A reasonable definition for this would be that, for the sequence of $T$-local observables $(K_{N})$, and for any real $\tau$, we have
@@
On the lower bound on the resolvent in (1.5). The lower bound $\Im R^{z}(e,e)\geq C_{0}^{-1}$ in (1.5) is only needed in the proofs of small scale quantum weak mixing to guarantee that for some $C^{\prime}_{0}>0$, for any $\eta>0$, all $N$ large enough and all $E\in I$, we have $|\Lambda_{J^{\eta}_{E}}|\geq C^{\prime}_{0}N\eta$. If one is interested in quantum ergodicity (1.8) or quantum weak mixing (1.18) on a fixed interval $I\subset I_{0}$, then a lower bound of the form $|\Lambda_{I}|\geq C^{\prime}_{0}N$ is sufficient (where $C^{\prime}_{0}$ is allowed to depend on $I$). The latter condition is satisfied under the sole hypothesis that the spectrum of the operator $P=\lambda(p)$ intersects the interior of $I$. See Remark 3.3 for more details.
@@ 2 Quantum mixing and resolvent estimates
2.1 Small scale quantum weak mixing implies quantum ergodicity and quantum weak mixing
@@
with the convention that $L_{I,J}=L_{I}^{\tau,\eta}=0$ if $\Lambda_{I}$ is empty. The notation is slightly abusive since $L^{\tau,\eta}_{I}(A,K)$ actually depends on the choice of the orthonormal basis. The quantum moment $L_{IJ}(A,K)$ is however independent of the choice of the basis as it can be written in terms of spectral projections (see the proof of Lemma 2.2 below). Note that, for centered observables $K$, the term $L^{0,0}_{I}$ appears in the definition of quantum ergodicity (1.8), the term $L^{\tau,\eta}_{I}$ in the definition of quantum weak mixing (1.18), while $L_{IJ}$ appears in the definition of small scale quantum weak mixing (1.11). We will be interested in situations where these moments are small. Since $L_{IJ}(A,K)$ is unitarily invariant, it will be easier to upper bound. However for $|\Lambda_{J}|$ large, $L_{IJ}(A,K)$ cannot be expected to be small in general.
Our strategy to upper bound $L_{I}^{\tau,\eta}$ will be to decompose $I$ into small sets counting few eigenvalues. The following lemma shows that small scale quantum weak mixing implies quantum ergodicity and quantum weak mixing (using the second part of the statement with $K_{N}-\langle K_{N}\rangle$ instead of $K_{N}$).
Lemma 2.1 (Small scale quantum weak mixing implies quantum ergodicity and mixing).
@@ 2.2 A general bound on quantum variance
The quantity $L_{IJ}(A,K)$ introduced in (2.1) is the central object that must be bounded from above in order to prove small scale quantum weak mixing. One key tool to control this quantity is to observe that it can be expressed as a trace involving $K$ and two functions of the operator $A$.
@@ 2.3 Proof ideas
$\left|\{g\in\Gamma\mid x=g.y\text{ and }f(p)_{g}\neq 0\}\right|>1.$
@@
$|f(P_{N})(x,y)|^{2}=\Bigl|\sum_{g\in\Gamma,\ x=g.y}f(p)_{g}\Bigr|^{2}=\sum_{g\in\Gamma,\ x=g.y}|f(p)_{g}|^{2}=\rho_{N}(q)(x,y)$ (2.6)
@@ 2.5 Resolvent identities
We explained in the previous section that the imaginary part of the resolvent appears in the trace computations needed to prove our main results. We gather here some elementary facts regarding resolvents. Let $X$ be a countable set and $P$ be a bounded self-adjoint operator on $\ell^{2}(X)$. In our case, we are interested in $X=\Gamma$, $P=\lambda(p)$ with $p=p^{*}\in\mathbb{C}[\Gamma]$. For $z\in\mathbb{C}^{+}=\{z\in\mathbb{C}:\Im(z)>0\}$, we define the resolvent of $P$ as $R^{z}=(P-z\mathrm{Id})^{-1}$. We will make use of the following resolvent identity, known as Ward’s identity: for any $x\in X$,
@@
where $\eta=\Im(z)$. It is a consequence of the fact that $\sum_{y\in X}|R^{z}(x,y)|^{2}=(R^{z}(R^{z})^{*})(x,x)$ together with the spectral theorem which implies that $\eta R^{z}(R^{z})^{*}=\Im R$ since $\Im(1/z)=-\Im(z)/|z|^{2}$ and $\Im(P-z\mathrm{Id})=-\eta\mathrm{Id}$, $|P-z\mathrm{Id}|^{-2}=R^{z}(R^{z})^{*}$. In the same vein, writing $z=E+i\eta$ and denoting by $\mu_{P}^{\delta_{x}}$ the spectral measure at vector $\delta_{x}$ (see (3.6)), we have
@@ 3.2 The number of eigenvalues in small intervals
We next take as usual $\rho_{N}$ converging in distribution toward $\lambda$ as $N\to\infty$. For an interval $J\subset I$, our goal in this subsection is to compare the number of eigenvalues of $\rho_{N}(p)$ in $J$ and $N\mu_{p}(J)$. For this, we prove the weak convergence to $\mu_{p}$ of the empirical distribution of eigenvalues of $\rho_{N}(p)$. More precisely, we establish the following uniform result.
@@ 3.3 Trace estimate for local operators
$A(x,y)=\sum_{g\in T}a_{g}(x){\mathbf{1}}(x=g.y)$ (3.11)
where $a_{g}\in\mathbb{C}^{N}$. Such a decomposition always exists: if for fixed $x,y$ there is a unique $g$ such that $x=g.y$, then $a_{g}(x)$ is uniquely defined by $a_{g}(x):=A(x,g^{-1}.x)$. This will be the typical situation considered below. If not, the decomposition is not unique. One could for example pick arbitrarily one $g$ such that $x=g.y$, put $a_{g}(x)=A(x,g^{-1}.x)$ and $a_{t}(x)=0$ for $t\neq g$ such that $t^{-1}.x=g^{-1}.x=y$. In the sequel, we make the more canonical choice $a_{g}(x)=A(x,g^{-1}.x)/m_{g}(x)$ where $m_{g}(x):=\sum_{t\in T}\mathbf{1}(g^{-1}.x=t^{-1}.x)\geq 1$.
@@
The goal of this section is to prove Corollary 3.5, which will be used in conjunction with Lemma 2.2. The result, roughly speaking, rephrases the trace in a dynamical way, as a scalar product of the observables $\langle K_{2},Q_{N}K_{1}\rangle$, where $Q_{N}$ is a propagator living on the finite graph. From there, we will either use Cauchy-Schwarz and relate $Q_{N}$ to a limiting operator $Q$ with nice decay properties using the convergence assumptions (Theorems 1.5-1.7), or use more specific properties of $K_{i}$ to conclude (Theorem 1.3). We need our propagator $Q_{N}$ to have the form $\rho_{N}(q)$ for some $q$ to be able to apply our assumptions, and this will hold true up to Benjamini-Schramm errors, which are controlled thanks to the assumption of convergence in distribution. In the case of diagonal observables $K_{1},K_{2}$ (corresponding to $T_{1}=T_{2}=\{e\}$ below), the statement simplifies dramatically and boils down to (2.5)-(2.6)-(2.7).
@@ 4.1 Quantum mixing estimate for Schreier graphs
We start by choosing $f_{j}$ and give some of its properties. In the sequel, we assume $\eta\leq 1/4$, which is no restriction since in the end we are interested in the limit $\eta\rightarrow 0$. We set
$\epsilon=\frac{1}{\ln\left(\frac{1}{\eta}\right)}$ (4.3)
@@
$\leq 4\sqrt{\int\frac{\eta^{2}}{(\lambda-E_{j})^{2}+\eta^{2}}d\mu_{p}(\lambda)}+4\eta$
@@
We may now evaluate (4.2) for our choice of $f_{j}$. We denote by $S$ a finite symmetric generating set of $\Gamma$ containing the support of $p$ and $t$, where we recall $t\in\Gamma$ is as in (4.1). For integer $r\geq 1$, $B_{S}(r)$ denotes the ball of radius $r$ in $\mathrm{Cay}(\Gamma,S)$ and ${\rm Bad}_{S}(r)=\mathrm{Fix}^{*}_{N}(B_{S}(2r))$. From Corollary 3.5, we have
@@
Since $f_{i}\geq 0$, we have $q=a^{*}a$ for some $a\in\mathbb{C}[\Gamma]$ as a consequence of Schur product theorem in $\mathbb{C}[\Gamma]$ and thus $Q_{N}=\rho_{N}(q)=\rho_{N}(a)^{*}\rho_{N}(a)$ is non-negative. From (4.4)-(4.5) and $\|k_{N}\|_{\infty}\leq 1$, we deduce our main upper bound: for some new $C>0$,
@@
Depending on the various assumptions stated in Theorems 1.3, 1.5 and 1.7, we will prove upper bounds on $\frac{\langle k_{N},Q_{N}k_{N}\rangle}{N\eta}$ which will imply (4.9).
@@
where we have used the bound $|q_{g}|=|f_{1}(p)_{t^{-1}gt}|\,|f_{2}(p)_{g}|\leq C_{1}^{2}\eta^{2}$ from (4.6). Note that this upper bound is uniform over all $E_{1},E_{2}\in I$. We deduce that
@@
Since $\delta$ can be arbitrarily small, this proves (4.9) and concludes the proof of Theorem 1.3.∎